### v1

@@ 9 Conclusions and outlook
Here, we presented a numerical framework to simulate moving boundary problems in biophysics. This work addresses a long-standing need for robust computational frameworks to tackle such problems. While other frameworks exist [85, 6, 7, 83, 13], our work targets flexibility of application while maintaining accuracy. Algorithms that enjoy stability and convergence properties are tied to non-invasive postprocessing techniques so to favor complex applications. The result is a stable pipeline with the capability of being scaled to more realistic scenarios.
Specifically, structure preservation is introduced on moving bulk and surface domains as a mean to allow for biophysically justified models and guarantee interpretability. We show that the scheme is readily implementable while being fast and accurate. The flexibility of the algorithm, which does not require ad hoc implementation, is proven by applying it to both advection-dominant ADR equations and phase-field models of the Cahn-Hilliard type. Convergence studies were performed for the bound- and mass-preserving scheme and the expected convergence rates were numerically verified. The framework also couples recent mesh redistribution techniques for gradient flows with the ALE method. We showed that this union is efficient in handling complex reshaping dynamics without the need for remeshing. As a result, the algorithm is tunable to the needs of the user and does not modify the underlying domain evolution. This is crucial in biophysical settings where the dynamics of the overall system is strongly coupled to the domains’ shape. Convergence studies were performed, demonstrating the accuracy of the scheme. Additionally, complex benchmark tests were conducted, showing good agreement with previous results in the literature. Building on previous results, we show that staggering is an effective method to simulate increasingly complex physics such as those that occur in biophysics.
The framework developed here has strong potential to simulate problems with increasing biological complexity. To achieve this goal, we foresee the need to incorporate large systems of ADR equations, that are not discussed in the present work [50, 83, 82]. Furthermore, fluid equations have also been shown to play an important role at certain timescales [94, 96, 2]. Including Navier-Stokes type equations on both interfaces and bulk [85, 7] will broaden the applicability of the computational schemes presented here.
@@ Acknowledgments
We acknowledge Dr. Emmet Francis for his valuable input on biologically relevant models for computational biology and for sharing the lessons learned from his programming experience [50].

### v2

@@ 9 Conclusions and outlook
Here, we presented a numerical framework to simulate moving boundary problems in biophysics. This work addresses a long-standing need for robust computational frameworks to tackle such problems. While other frameworks exist Mokbel et al. [2024], Arroyo and DeSimone [2009], Bachini et al. [2023], Mackenzie et al. [2021], Barrett et al. [2017b], our work targets flexibility of application while maintaining accuracy. Algorithms that enjoy stability and convergence properties are tied to non-invasive postprocessing techniques so to favor complex applications. The result is a stable pipeline with the capability of being scaled to more realistic scenarios.
Specifically, structure preservation is introduced on moving bulk and surface domains as a mean to allow for biophysically justified models and guarantee interpretability. We show that the scheme is readily implementable while being fast and accurate. The flexibility of the algorithm, which does not require ad hoc implementation, is proven by applying it to both advection-dominant ADR equations and phase-field models of the Cahn-Hilliard type. Convergence studies were performed for the bound- and mass-preserving scheme and the expected convergence rates were numerically verified. The framework also couples recent mesh redistribution techniques for gradient flows with the ALE method. This union is showed to be efficient in handling complex reshaping dynamics without the need for remeshing. As a result, the algorithm is tunable to the needs of the user and does not modify the underlying domain evolution. This is crucial in biophysical settings where the dynamics of the overall system is strongly coupled to the domains’ shape. Convergence studies were performed, demonstrating the accuracy of the scheme. Additionally, complex benchmark tests were conducted, showing good agreement with previous results in the literature. Building on previous results, we show that staggering is an effective method to simulate increasingly complex physics such as those that occur in biophysics.
The framework developed here has strong potential to simulate problems with increasing biological complexity. To achieve this goal, we foresee the need to incorporate large systems of ADR equations, that are not discussed in the present work Francis et al. [2024], Mackenzie et al. [2021], MacDonald et al. [2016]. Furthermore, fluid equations have also been shown to play an important role at certain timescales Seifert [1997], Shin and Brangwynne [2017], Alberti and Hyman [2021]. Including Navier-Stokes type equations on both interfaces and bulk Mokbel et al. [2024], Bachini et al. [2023] will broaden the applicability of the computational schemes presented here.
@@ Acknowledgments
We acknowledge Dr. Emmet Francis for his valuable input on biologically relevant models for computational biology and for sharing the lessons learned from his programming experience Francis et al. [2024]. All the simulations are performed using the software NgSolve Schöberl [2014], Gangl et al. [2021].